OpenStudy (jcmbourque19):

Someone PLEASE HELP!!! I've been stuck for HOURS. I will work with you, I don't just want the answer, I need to know how and why. PLEASE!!! What is a40 for the arithmetic sequence presented in the table below? n 8 12 an 60 48 Hint: an = a1 + d(n − 1), where a1 is the first term and d is the common difference.

OpenStudy (peachpi):

you can think of an arithmetic sequence as a line. So then (n, an) would be a point on the line. So start by finding the equation of the line that goes through (8, 60) and (12, 48). Do you know how to do that?

OpenStudy (jcmbourque19):

The slope is the d value right? So would it be -3?

OpenStudy (peachpi):

yes

OpenStudy (jcmbourque19):

Okay so what now? Is a1 8?

OpenStudy (peachpi):

They didn't actually give you a1. You can actually start this in point slope form, then back out an equation for the sequence y -y1 = m(x-x1) or with your terms an - a8 = d(n - 8) the 12 can replace the 8 just as easily

OpenStudy (peachpi):

does that make sense?

OpenStudy (jcmbourque19):

It makes sense but I've only been given one formula, the one listed above. This is the first lesson of the chapter, so I'd like to solve it the way they want me to. I really appreciate your help, btw. I'm just so stuck. I've done other problems using this formula successfully and I don't know why this question is so confusing.

OpenStudy (peachpi):

this one can't be solved with that formula alone because that formula requires you to have the 1st term of the sequence. You were given the 8th and 12th terms. So you have to either write the sequence using the 8th and 12th terms (above) or back out the 1st term by subtracting -3 from the 8th and subsequent terms

OpenStudy (jcmbourque19):

What would be the value of n? My answer choices for a40 are -44, -40, -36, and -26.

OpenStudy (peachpi):

ok maybe let's try it this way. We know the 8th term is 60 and d is -3, so you can solve for the 1st term \[a_8=a_1+d(n-1)\] \[60=a_1-3(8-1)\] Solve that for \(a_1\)

OpenStudy (jcmbourque19):

So the first term is 81 then. I worked backwards and figured that out previously, but I still don't know what to do for n.

OpenStudy (peachpi):

n is 40

OpenStudy (peachpi):

because they ask for a40

OpenStudy (peachpi):

so then you can you your formula \[a_n=a_1+d(n-1)\] \[a_{40}=81-3(40-1)\]

OpenStudy (jcmbourque19):

so a40=81-3(40-1) a40=81-117 a40=-36

OpenStudy (peachpi):

yes

OpenStudy (jcmbourque19):

For like three hours I've been adding in an extra - sign and getting 198. Thank you so much for your help!!! How do I give you a medal?

OpenStudy (peachpi):

you're welcome. seems like you figured it out :)

OpenStudy (jcmbourque19):

If I have any questions in the future, would you mind if I tag you?